Let R={(1,2),(2,1),(2,3),(3,4),(4,1) } be a relation on the A={1,2,3,4}. Find the transitive closure of R using Warshall’s algorithm.


1
Expert's answer
2021-12-21T18:34:32-0500

Consider a relation "R=\{ (1,2),(2,1),(2,3),(3,4),(4,1)\}" on the set "A=\{1,2,3,4\}".


Let us state the steps of the Warshall's algorithm:


Step 1. Let "W:=M_R,\ k:=0."

Step 2. Put "k:=k+1."

Step 3. For all "i\ne k" such that "w_{ik}=1" and for all "j" let "w_{ij}=w_{ij}\lor w_{kj}."

Step 4. If "k=n" then STOP: "W=M_{R^*}," else go to Step 2.


Let us find transitive closure of the relation "R" using Warshall's algorithm:


"W^{(0)}=M_R
=\begin{pmatrix}
0 & 1 & 0 & 0\\
1 & 0 & 1 & 0\\
0 & 0 & 0 & 1\\
1 & 0 & 0 & 0
\end{pmatrix}"


"W^{(1)}
=\begin{pmatrix}
0 & 1 & 0 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
1 & 1 & 0 & 0
\end{pmatrix}"


"W^{(2)}
=\begin{pmatrix}
1 & 1 & 1 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 0
\end{pmatrix}"


"W^{(3)}
=\begin{pmatrix}
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


"M_{R^*}=W^{(4)}
=\begin{pmatrix}
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


It follows that "R^*=A\times A" is a universal relation on the set "A."